OpenStudy (anonymous):

solve the equation by finding square roots 3x^2=21

jhonyy9 (jhonyy9):

divide both sides by 3

OpenStudy (diyadiya):

\[\sqrt{3x^2} = \sqrt{21}\]Now divide both sides by \(\sqrt{3}\)\[\sqrt{x^2} = \frac{\sqrt{21}}{\sqrt{3}}\]\[x= \frac{\sqrt{21}}{\sqrt{3}}\]\[x= \frac{\sqrt{7 } \times \sqrt{3}}{\sqrt{3}} = \frac{\sqrt{7 } \times \cancel{\sqrt{3}}}{\cancel{\sqrt{3}}}\] so x= ?

OpenStudy (anonymous):